Recent Log Entries

FSD Masking Tool: Successful System UN-Masking for: CryoModule

Log Number: 4186944Logbook(s): Tags: Body: System UN-Masking for: CryoModule
User: jhoward
Areas: All

CryoModule Devices:
0L02
0L04
0L03
1L14
1L03
1L02
1L04
1L05
1L07
1L06
1L08
1L10
1L11
1L09
1L12
1L15
1L13
1L16
1L19
1L17
1L18
1L20
1L21
1L22
1L23
1L24
1L25
1L26
2L14
2L03
2L02
2L04
2L07
2L05
2L06
2L08
2L11
2L09
2L10
2L12
2L15
2L13
2L16
2L19
2L17
2L18
2L20
2L21
2L22
2L23
2L24
2L25
2L26


QE Measurement

Log Number: 4186942Logbook(s): Body: Spot Location: (12590.0,18700.0), Charge: 18.833174999999997

Laser Wavelength Atten Setting Laser Power Current HV PS Current QE PV Updated? Update Time/Reason Not Updated (nm) (W) (uA)</cente

MQR1C20 Cycled (Resistance Alarm)

Log Number: 4186943Logbook(s): Entry Makers: emonkBody: 

To clear the resistance alarm, MQR1C20 of Hall A was cycled.

CW Override Switch was enabled.

Figure 1: MQR1C20 cycling graph.

FSD Masking Tool: Successful System Masking for: CryoModule

Log Number: 4186936Logbook(s): Tags: Body: System Masking for: CryoModule
User: jhoward
Areas: All

CryoModule Devices:
0L02
0L04
0L03
1L14
1L03
1L02
1L04
1L05
1L07
1L06
1L08
1L10
1L11
1L09
1L12
1L15
1L13
1L16
1L19
1L17
1L18
1L20
1L21
1L22
1L23
1L24
1L25
1L26
2L14
2L03
2L02
2L04
2L07
2L05
2L06
2L08
2L11
2L09
2L10
2L12
2L15
2L13
2L16
2L19
2L17
2L18
2L20
2L21
2L22
2L23
2L24
2L25
2L26


COMMENT ON EIC Girder Unloading

Log Number: 4186930Logbook(s): Body: 

4 Girders delivered at WSMA, unloaded 9/20 & 2 more 9/21. On site girders awaiting purging then they will be moved into one of the tents for storage.

Hall A Target Change - 5uA of CW on PolHe3 cell

Log Number: 4186902Logbook(s): Entry Makers: emonkBody: 

Hall A requested - 5uA of CW on PolHe3 cell.

Ion Chambers were calibrated.
Max Juice remained the same (10uA).
XSOFs set.

Figure 1: iCalibrate Settings (5uA with 30% Margin for - PolHe3 cell)

Figure 2: Operational Restrictions

Figure 3: Hall A Target Verification Screen once raster, ion chambers, max juice, and XSOFs are set and target is in place

iCalibrate: Hall A

Log Number: 4186900Logbook(s): Body: Calibration ParametersHall:ATarget: Reference CellPass: 4Note: Refernce Cell/PolHe3Calibrated On:2023-Sep-15 12:15Calibrated By: schoeneFile:A-2023-09-15_1215_Reference_Cell_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

Autolog save

Log Number: 4186899Logbook(s): Tags: Body: saver ran from opsl03

Saved: emonk 09-25-23 02:55:07 {A:4 B:5 C:5 E:1047 A=CW B=BS C=CW D=BS 0R08=35.13 No dumps in.

Follow-up Re: ARC 1 BEM Losing Communcation

Log Number: 4186898Logbook(s): Body: 

The problem seems to be getting worse and it is affecting more than just ARC1 BEM. See attached figure. Looks like all the Arcs (minus ARC6) are experiencing this issue.

Additionally, the Hall C valid indicator is the only one that isn't archived.

Hall A Target Change to 5uA CW with No Target

Log Number: 4186892Logbook(s): Entry Makers: emonkBody: 

Hall A requested - 5uA of CW with no target.

Ion Chambers were calibrated.
Max Juice remained the same (15uA).
XSOFs set.

Figure 1: iCalibrate Settings (5uA with 30% Margin for - no target)

Figure 2: Operational Restrictions

Figure 3: Hall A Target Verification Screen once raster, ion chambers, max juice, and XSOFs are set and target is in place

iCalibrate: Hall A

Log Number: 4186891Logbook(s): Body: Calibration ParametersHall:ATarget: No TargetPass: 4Note: SBS and raster onCalibrated On:2023-Sep-25 01:23Calibrated By: moserFile:A-2023-09-25_0123_No_Target_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Compton</th

Downtime Incident CLOSED: BigBite Power Trip

Log Number: 4186890Logbook(s): Body: IDDescriptionPeriodCause12034Title:
BigBite Power Trip

Summary:
BigBite lost current this shifted positions enough to where we can't send beam without tripping Ion Chambers (specifically He3 Collimator).

Operator In Training SIGN IN

Log Number: 4186888Logbook(s): Body:  ACCELERATOR OPERATOR IN TRAINING SIGN IN FORM OPERATOR:    Monk
DATE:           09/25/2023
SHIFT:             Owl
MENTOR:         Merz

The operator has read the following:

Downtime Incident CREATED: BigBite Power Trip

Log Number: 4186887Logbook(s): Body: IDDescriptionPeriodCause12034Title:
BigBite Power Trip

Summary:
BigBite lost current this shifted positions enough to where we can't send beam without tripping Ion Chambers (specifically He3 Collimator).

Pages